# XTransfer: Cross-Modality Model Transfer for Human Sensing with Few Data at the Edge

### v1

@@ 4 Model Repairing
Enabled by SRR pipeline, this process safely repairs layer-wise misalignment by minimizing MMC shifts, effectively restoring layer performance from source models using few sensing inputs in cross-modality settings. SRR pipeline involves three key stages shown in Fig. 3. In the Splice stage, input shapes between source models and sensor data must be aligned well, as default reshaping (c.f. App. B.1)—using fixed down/up-sampling—cannot be fine-tuned and may lose important features across channel, height, and width, especially when layer shapes vary during layer-wise search. To enhance, we design a compact connector placed between sensing inputs and layers, consisting of a Pre-header (adaptive convolutional layer), Resizer, and one encoder-decoder pair to bridge cross-modality structures. In the next Repair stage, each connector is fine-tuned to generate adaptive channels, preserve key sensing features, and minimize MMC shift via the generative transfer module. In the following Removal stage, channel removal is designed to further enhance layer-wise performance while optimizing resource usage. For more design details, see Appendix B.2.
Figure 3: XTransfer overview. LWS control segments source models into layers and uses the pre-search check to decide if repairing is needed. SRR pipeline then fine-tunes connectors to repair selected layers. Finally, LWS control selects and recombines layers of interest into a compact model.
@@ 4.1 Repairing in complex feature space
Since the original latent features from source models are well-trained for high discriminability (c.f. Sec. 3), our intuition is to safely fine-tune the misinterpreted latent features of few sensor data by aligning with the original features (i.e., anchors) to minimize MMC shifts. We hence propose anchor-based cross-modality supervision alignment, which aligns source MMCs as anchors (i.e., anchor MMCs) with sensing MMCs. To handle complex feature spaces, our insight is to reduce MMC dimensionality by preserving key channels and excluding irrelevant ones.
@@
Reduced-dimensional feature space. We project anchor MMCs into a lower-dimensional feature space using principal component analysis (PCA) (VALPOLA2015143, ) (i.e., anchor PCA space). PCA enables a lower-dimensional subspace (i.e., adjusted by the number of PCA components, with 2 components used by default) where the projected data exhibits the highest possible variance (i.e., highly clustered). The optimized component weights (i.e., projection coefficients) highlight the most important channels that contribute to the performance of each layer. Given high accuracy achieved in each original layer, the projected anchor MMCs show high clustering performance, e.g., the centroid of the distribution of each class (i.e., Anchor-center) clustered tightly, as shown in Fig. 3(a). Once the anchor PCA space is initialized for each layer, we project sensing MMCs into its space using the same parameters.
@@
Feature space alignment. After projecting sensing MMCs, we observe that the sample distribution for each class, along with their centroid (i.e., Sensing-center), appears disorganized compared to the well-clustered Anchor-center at each layer, as shown in Fig. 3(a). Specifically, the scale and orientation of Sensing-center are notably different from those of Anchor-center, caused by MMC shifts. To align the scale, we first define a scale function based on the mean shift in inter-class distance $\mathrm{InterD}(\cdot)$, formulated as $S={{Mean}({InterD}(Pro(fs_{ij})))}\mathbin{/}{{Mean}({InterD}(Pro(ft_{ij})))}$, where $Pro(\cdot)={PCA}({MMC}(\cdot))$ denotes the space project function, $fs$ and $ft$ represent the latent features from both source and sensing modalities, respectively. $i$ denotes the index of pre-trained models, and $j$ denotes the layer index in each model $i$. Next, to align the orientation, we design a rotation alignment function that minimizes the cosine angle (nguyen2010cosine, ) by fine-tuning a multi-dimensional rotation matrix $[M_{rot}]$ (where the dimensions equal to the number of PCA components) as below:
$\footnotesize\begin{split}\arg\min_{[M_{rot}]}Cos(Cen(Pro(fs_{ij})),Cen(Pro(ft_{ij})*S)\cdot[M_{rot}])\end{split}$ (1)
@@
where $Cos(\cdot)$ denotes the cosine similarity function, and $Cen(\cdot)$ calculates the centroid of the projected distribution. Fig. 3(b) shows the aligned results.
@@
Anchor class pairing. Building on key finding that class pairing boosts latent feature alignment across modalities (c.f. Sec. 3), we propose selecting source classes with the highest S-scores as anchor classes to pair with sensing classes to facilitate the supervision alignment. We minimize pairing shift (i.e., the sum distance across each pair) via a linear sum assignment solved using the standard Hungarian algorithm, outputting the optimal pairing set $pairST$.
@@ 4.2 Repairing layer-wise dependence
Generative transfer module. To effectively restore layer-wise dependence, we design a generative transfer module that enhances the supervision alignment through generative learning. It consists of a generator (i.e., connector) and a discriminator—each frozen pre-trained layer paired with its anchor PCA space (Fig. 3). Freezing preserves stable anchors and prevents overfitting. For each layer, given the optimal pairing set $pairST$, the generator is fine-tuned using the discriminator’s loss to minimize MMC shifts, and uniquely transforms sensor inputs to align with anchors.
Anchor-based repair loss. Building on the generative transfer module, we propose an anchor-based repair loss function ($loss_{srr}$) to harmonize latent feature distributions across modalities. Given each layer’s anchor PCA space, our objective is twofold: 1) minimize the Euclidean distance (ED) ($D(\cdot)$) between the centroid of projected anchor MMC distribution and that of the projected sensing MMC distribution for each class pair $(c_{s},c_{t})\in pairST$ (i.e., positive loss), and 2) maximize the distance between sensing MMC samples with different labels $(c_{t}^{-})$ (i.e., negative loss). Since the projected anchor MMCs are highly clustered, it can provide the maximum margin ($M_{max}$) between anchor clusters. To further improve the negative loss, we update the function by calculating the anchor-based margin $margin^{c}=InterD(Pro(fs^{c}_{ij}))-IntraD(Pro(fs^{c}_{ij}))$, where the $M_{max}$ is selected, and $IntraD(\cdot)$ measures the intra-class distance. Hence, our loss function is formulated as:
$\footnotesize\begin{split}loss_{srr}=&\frac{1}{N}\sum^{N}_{(c_{s},c_{t})\in pairST}D(Cen(Pro(fs^{c_{s}}_{ij})),Cen(Pro(ft^{c_{t}}_{ij})))\\ +&\frac{1}{K}\sum^{K}_{(c_{t},c_{tk}^{-})}ReLU(M_{max}-D(Pro(ft^{c_{t}}_{ij}),Pro(ft^{c_{tk}^{-}}_{ij})))\end{split}$ (2)
where $K$ denotes the number of negative samples, and $N$ denotes the number of classes. The $ReLU(\cdot)$ is used to trigger the negative loss in the condition where the distance between the samples is less than $M_{max}$. Fig. 3(c) shows the effectiveness of model repairing.
@@ 5 Layer Recombining
Working in conjunction with SRR pipeline for repairing each layer, XTransfer leverages layer recombining enabled by LWS control to: 1) perform the pre-search check that accelerates the search process, and 2) efficiently select and recombine layers of interest to form a compact model.
@@ 5.1 Search control design
Following NAS principles (searchinsights, ), poor design choices (e.g., search space, actions, strategy) can disrupt the balance between model performance and resource efficiency. As layers of interest grow, resource overhead can increase exponentially, making dynamic search control essential under on-device resource constraints. To efficiently manage, we define the search space as all layers from source models, structured into search pools containing layer candidates (i.e., repaired layers), each with width $I$ (number of models) and depth $J$ (number of L-units), forming a window of size $I\times J$. Search actions (init, continue, skip, cross) are defined to guide layer recombination within each designated search pool. Init selects the starting layer with a default fixed depth of 1. Continue and skip guide intra-model layer recombination, while cross enables inter-model layer recombination.
Search strategy under resource constraints. To effectively control search actions within each search pool, we propose a layer value function based on the S-score $Score(\cdot)$ to evaluate both individual repaired layer performance and the convergence between adjacent layers after each SRR process $SRR(\cdot)$. It is formulated as $V_{ij}=Score(SRR(Pro(ft_{ij}),MT_{ij},L_{ij}),Y_{t}))$, where $MT_{ij}$ denotes the connector and $L_{ij}$ represents each repaired layer. Under resource constraints, only the top-performing layer in each pool is selected. LWS control then checks whether its S-score exceeds that of the previous layer. If successful, the layer $L_{ij}$ is identified as the layer of interest and recombined, and the next pool window begins at layer index $j+1$. Otherwise, the current pool is discarded, and the search window shifts forward. Considering the constraints, we design a resource-constrained function that works in conjunction with $V_{ij}(\cdot)$, which is formulated as $RC(n)=exp(n/(L-2)-2)+1,n\in[1,~{}L]$, where $n$ denotes the current recombined layer index, and $L$ denotes the maximum number of L-units. Each layer’s adjusted overhead, $R(n)_{ij}$, is computed as $R(n)_{ij}=(Res(L_{ij})*RC(n))/Max(Res(P))$, where $Res(\cdot)$ calculates actual cost (e.g., FLOPs), and $Max(Res(P))$ is the peak cost in current pool $P$. The final resource-constrained layer value function becomes $VR_{ij}=V_{ij}\mathbin{/}R(n)_{ij}$. The objective is then defined as $\arg\max_{L_{ij}\in P}VR_{ij},\,s.t.\sum_{0}^{n}Res(L_{ij})\leq R_{Max}$, where $R_{Max}$ is the total allowed resource budget.
@@ 5.2 Efficient search
Pre-search check. LWS control is designed to handle large search spaces from deep or multi-source models. While fine-tuning the connectors in SRR pipeline is efficient, the large search space can considerably slow down LWS control. We hence propose the pre-search check strategy to avoid unnecessary repair steps before applying SRR. We track S-score progression across three states—after anchor pairing (Anchor) (i.e., repairing objective), before SRR (Before), and after SRR (After) (i.e., repaired). SRR pipeline consistently improves S-scores from Before to After, approaching the Anchor level, indicating a strong correlation. Based on this, our intuition is to estimate repaired S-score (i.e., repair rate) of each layer based on Before and Anchor states to decide whether to bypass unnecessary repairs. Experiments show the actual repair rate grows exponentially with the recombined layer index (see App. B.4). We hence design a repair rate growth model formulated as $rate_{n}^{est}=exp(a*n)+b,n\in[1,L]$, where $a$ and $b$ are optimized via non-linear regression to minimize MSE to fit $rate_{n}^{est}$, with the objective function defined as $\arg\min_{a,b}\frac{1}{n^{{}^{\prime}}}\sum_{n=0}^{n^{{}^{\prime}}}(rate_{n}^{est}-rate_{n})^{2}$, where $n^{{}^{\prime}}$ is the number of recombined layers. The estimation begins after init. Once a new layer is selected, LWS control updates the estimation. Due to the limited updates, the estimation may be inaccurate. To mitigate, we also develop a dynamic search range mechanism that dynamically adjusts the search range in each pool (details in App. B.5).
@@ 6 Experiment
6.1 Experiment setup
Table 1: State-of-the-art baselines.
@@ 6.1 Experiment setup
  TL (TransferabilityID, ) Single-source A standard transfer learning method by freezing all layers
FT (Jaehoon2022, ) Single-source A standard fine-tuning method by updating all layers
ProtoNet (Snell2017, ) Single-source A distance-based FSL method
MAML (Finn2017, ) Single-source A learning-based FSL method
DAPN (Zhao2021, ) Single-source A cross-domain FSL method using adversarial ProtoNet
GPT2 (zhou2023one, ) Single-(LLM) A large language model based method using fine-tuning
MDDA (Sicheng2020, ) Multi-source A cross-domain FSL method using knowledge distillation
MCW (NEURIPS20196048ff4e, ) Multi-source A cross-domain FSL method using maximal correlation analysis
MetaSense (metasense, ) Oracle-(GT) An advanced FSL method based on MAML with sufficient target data
@@
miniImageNet (miniimagenet, ) Image - 100 - 3x84x84 Public
miniDomainNet (domainnet, ) Image - 126 - 3x224x224 Public
Office-31 (office31, ) Image - 31 - 3x224x224 Public
Office-Home (officehome, ) Image - 65 - 3x224x224 Public
Caltech-101 (caltech101, ) Image - 101 - 3x224x224 Public
Newsgroup (text2021, ) Text - 20 - 100x1x1000 Public
VoxCeleb (VoxCeleb2017, ) Audio - 100 Time-S 1x512x300 Public
OPPORTUNITY (Roggen2010, ) IMU 12 11 Time-S 77x1x100 Public
@@
HHAR (Stisen2015, ) IMU 9 6 Time-S 6x1x256 Public
WESAD (Schmidt2018, ) ECG 15 3 Time-S 10x1x256 Public
@@
ChestX (Wang2017, ) Image - 5 - 1x1024x1024 Public
@@
  ResNet18 (chen2018a, ) 18 9 3.67 11.18M 2D 3x224x224
ResNet18-1D (chen2018a, ) 18 9 0.39 3.9M 1D 100x1x1000
Multi-ResNet18 (Sicheng2020, ) 18x5 9 18.35 55.9M 2D 3x224x224
Conv4 (chen2018a, ) 4 4 0.02 0.07 1D 1x1x500
Conv5 (metasense, ) 5 5 0.19 30.42M 1D By Input
GPT2 (zhou2023one, ) 4x6 6 0.41 82.9M 1D By Input
@@
We benchmark XTransfer against state-of-the-art baselines, grouped into single- and multi-source methods (Tab. 1). Evaluation is conducted on 8 widely-used public datasets for cross-modality model transfer (Tab. 2), using their corresponding public pre-trained models as sources, from diverse modalities (e.g., image, text, audio, and sensing) with standard backbones such as ResNet18 from PyTorch Hub (PyTorchhub, ) (Tab. 3). To evaluate XTransfer, we develop 4 real-world applications and build three testbeds on 3 commercial edge devices (e.g., smartwatch, smartphone, raspberry pi). Based on these, we collect 4 human sensing datasets (marked as Private in Tab. 2) as targets, from 40 participants in real-world settings, with ethics approval from the Human Research Ethics Committee of our institute. Additionally, we include 2 public sensing datasets and 1 image dataset as targets for standard benchmarking. We apply standard Leave-One-Out Cross-Validation (LOOCV) (metasense, ) for evaluating human sensing datasets, and adopt accuracy-to-resource (ATR) ratio, defined as $ATR=Accuracy\mathbin{/}(\alpha Norm(FLOPs))+(1-\alpha)Norm(Params))$ for evaluating resource-accuracy efficiency, where $\alpha$ is set to 0.5 by default. For more details, see Appendix A.

### v2

@@ 4 Model Repairing
Enabled by SRR pipeline, this process safely repairs layer-wise misalignment by minimizing MMC shifts, effectively restoring layer performance from source models using few sensing inputs under cross-modality FSL settings. Three stages are designed, as shown in Figure 1. In the Splice stage, default reshaping (see Appendix B.1)—using fixed down/up-sampling—cannot be fine-tuned and may lose important features, especially when layer shapes vary during layer-wise search. We propose a compact connector placed between sensing inputs and layers, consisting of a Pre-header (adaptive convolutional layer), Resizer, and one encoder-decoder pair to bridge cross-modality structural gaps. In the Repair stage, each connector is fine-tuned by the generative transfer module to generate adaptive channels, preserve key sensing features, and minimize MMC shifts. Finally, the Removal stage applies channel removal (see Appendix B.2) to further enhance layer-wise performance.
@@ 4.1 Repairing in complex feature space
Since the original latent features from source models are well-trained for high discriminability, our intuition is to safely fine-tune few sensor data by aligning the misinterpreted latent features with the original features (i.e., anchors) to minimize MMC shifts. We hence propose cross-modality anchor-based alignment, which aligns source MMCs as anchors (i.e., anchor MMCs) with sensing MMCs. To handle complex feature spaces, our insight is to reduce MMC dimensionality by preserving key channels and excluding irrelevant or noisy ones.
@@
Reduced-orthogonal feature space. We project anchor MMCs into a lower-dimensional, orthogonal feature space using principal component analysis (PCA) (Valpola, 2015) (i.e., anchor PCA space). Unlike the raw MMC feature space, where channel dimensions are often noisy and correlated, PCA enforces orthogonality across components, removing redundancy and suppressing noise. This enables a compact subspace (i.e., defaulting to 2 components) where the projected data captures maximum variance and forms highly clustered distributions. The optimized component weights (i.e., projection coefficients) highlight the most important channels contributing to layer performance. Given high discriminability of the original layers, the projected anchor MMCs preserve high clustering performance, with class distributions tightly centered (i.e., Anchor-center), as shown in Figure 1(a). Once initialized for each layer, this anchor PCA space is reused to project sensing MMCs.
@@
Feature space alignment. After projecting sensing MMCs, we observe that the sample distribution for each class, along with their centroid (i.e., Sensing-center), appears disorganized compared to the well-clustered Anchor-center at each layer, as shown in Figure 1(a). Specifically, the scale and orientation of Sensing-center are notably different from those of Anchor-center, caused by MMC shifts. To align the scale, we first define a scale function based on the mean shift in inter-class distance $\mathrm{InterD}(\cdot)$, formulated as $S={{Mean}({InterD}(Pro(fs_{ij})))}\mathbin{/}{{Mean}({InterD}(Pro(ft_{ij})))}$, where $Pro(\cdot)={PCA}({MMC}(\cdot))$ denotes the space project function, $fs$ and $ft$ represent the latent features from both source and sensing modalities, respectively. $i$ denotes the index of pre-trained models, and $j$ denotes the layer index in each model $i$. Next, to align the orientation, we design a rotation alignment function that minimizes the cosine angle (Nguyen & Bai, 2010) by fine-tuning a multi-dimensional rotation matrix $M_{rot}]$ with dimensions matching the number of PCA components:
$\begin{split}\arg\min_{[M_{rot}]}Cos(Cen(Pro(fs_{ij})),Cen(Pro(ft_{ij})*S)\cdot[M_{rot}])\end{split}$ (1)
@@
where $Cos(\cdot)$ denotes the cosine similarity function, and $Cen(\cdot)$ calculates the centroid of the projected distribution. Figure 1(b) shows the aligned results.
@@
Anchor class pairing. Building on the key finding that class pairing boosts latent feature alignment across modalities (see Section 3), we propose selecting source classes with the highest S-scores as anchor classes to pair with sensing classes to facilitate the anchor-based alignment. We minimize pairing shift (i.e., the sum distance across each pair) via a linear sum assignment solved using the standard Hungarian algorithm, outputting the optimal pairing set $pairST$.
@@ 4.2 Repairing layer-wise dependence
Generative transfer module. To effectively restore layer-wise dependence, we design a generative transfer module that enhances the anchor-based alignment through generative learning. It consists of a generator (i.e., connector) and a discriminator—each frozen pre-trained layer paired with its anchor PCA space (Figure 1). Freezing preserves stable anchors and prevents overfitting. For each layer, given the optimal pairing set $pairST$, the generator is fine-tuned using the discriminator’s loss to minimize MMC shifts, and uniquely transforms sensor inputs to align with anchors.
Anchor-based repair loss. Building on the generative transfer module, we propose an anchor-based repair loss function ($loss_{srr}$) to harmonize latent feature distributions across modalities. Given each layer’s anchor PCA space, our objective is twofold: 1) minimize the distance ($D(\cdot)$) between the centroid of projected anchor MMC distribution and that of the projected sensing MMC distribution for each class pair $(c_{s},c_{t})\in pairST$ (i.e., positive loss), and 2) maximize the distance between sensing MMC samples with different labels $(c_{t}^{-})$ (i.e., negative loss). Since the projected anchor MMCs are highly clustered, they can provide the maximum margin ($M_{max}$) between anchor clusters. To further improve the negative loss, we update the function by calculating the anchor-based margin $margin^{c}=InterD(Pro(fs^{c}_{ij}))-IntraD(Pro(fs^{c}_{ij}))$, where the $M_{max}$ is selected, and $IntraD(\cdot)$ measures the intra-class distance. Hence, our loss function is formulated as:
$\begin{split}loss_{srr}=&\frac{1}{N}\sum^{N}_{(c_{s},c_{t})\in pairST}D(Cen(Pro(fs^{c_{s}}_{ij})),Cen(Pro(ft^{c_{t}}_{ij})))\\ +&\frac{1}{K}\sum^{K}_{(c_{t},c_{tk}^{-})}ReLU(M_{max}-D(Pro(ft^{c_{t}}_{ij}),Pro(ft^{c_{tk}^{-}}_{ij})))\end{split}$ (2)
where $K$ denotes the number of negative samples, and $N$ denotes the number of classes. The $ReLU(\cdot)$ is used to trigger the negative loss in the condition where the distance between the samples is less than $M_{max}$. Figure 1(c) shows the effectiveness of model repairing.
@@ 5 Layer Recombining
Working in conjunction with SRR pipeline for repairing each layer, XTransfer leverages layer recombining enabled by LWS control to: 1) perform the pre-search check that accelerates the search process, and 2) effectively select and recombine layers of interest to form a compact model.
@@ 5.1 Search control design
Following NAS principles (White et al., 2023), poor design choices (e.g., search space, actions, strategy) can disrupt the balance between model performance and resource efficiency. As layers of interest grow, resource overhead can increase exponentially, making dynamic search control essential under on-device resource constraints. To effectively manage, we define the search space as all layers from source models, structured into search pools containing layer candidates (i.e., repaired layers), each with width $I$ (number of models) and depth $J$ (number of L-units), forming a window of size $I\times J$. Search actions (init, continue, skip, cross) are defined to guide layer recombination within each designated search pool. Init selects the starting layer with a default fixed depth of 1. Continue and skip guide intra-model layer recombination, while cross enables inter-model layer recombination.
Search strategy under resource constraints. To effectively control search actions within each search pool, we propose a layer value function based on the S-score $Score(\cdot)$ to evaluate each repaired layer performance and the convergence between adjacent layers after each SRR process $SRR(\cdot)$. It is formulated as $V_{ij}=Score(SRR(Pro(ft_{ij}),MT_{ij},L_{ij}),Y_{t}))$, where $MT_{ij}$ denotes the connector and $L_{ij}$ represents each repaired layer. Under resource constraints, only the highest-valued repaired layer in each pool is selected. LWS control then checks whether its S-score exceeds that of the previous layer to maintain layer-wise dependence. If successful, the layer $L_{ij}$ is identified as the layer of interest and recombined, and the next pool window begins at layer index $j+1$. Otherwise, the current pool is discarded, and the search window shifts forward. Considering the constraints, we design a resource-constrained function that works in conjunction with $V_{ij}(\cdot)$, which is formulated as $RC(n)=exp(n/(L-2)-2)+1,n\in[1,~L]$, where $n$ denotes the current recombined layer index, and $L$ denotes the maximum number of L-units. Each layer’s adjusted overhead, $R(n)_{ij}$, is computed as $R(n)_{ij}=(Res(L_{ij})*RC(n))/Max(Res(P))$, where $Res(\cdot)$ calculates actual cost (e.g., FLOPs), and $Max(Res(P))$ is the peak cost in current pool $P$. The resource-constrained layer value function becomes $VR_{ij}=V_{ij}\mathbin{/}R(n)_{ij}$. The objective is defined as $\arg\max_{L_{ij}\in P}VR_{ij},\,s.t.\sum_{0}^{n}Res(L_{ij})\leq R_{Max}$, where $R_{Max}$ is device-specific resource budget.
@@ 5.2 Efficient search
Pre-search check. LWS control is designed to handle large search spaces from deep or multi-source models. While fine-tuning the connectors in SRR pipeline is efficient, the large search space can considerably slow down LWS control. We hence propose the pre-search check strategy to avoid unnecessary repair steps before applying SRR. We track S-score progression across three states—after anchor pairing (Anchor) (i.e., repairing objective), before SRR (Before), and after SRR (After) (i.e., repaired). SRR pipeline consistently improves S-scores from Before to After, approaching the Anchor level, indicating a strong correlation. Based on this, our intuition is to estimate repaired S-score (i.e., repair rate) of each layer based on Before and Anchor states to decide whether to bypass unnecessary repairs. Experiments show the actual repair rate grows exponentially with the recombined layer index (see Appendix B.4). We hence design a repair rate growth model formulated as $rate_{n}^{est}=exp(a*n)+b,n\in[1,L]$, where $a$ and $b$ are optimized via non-linear regression to minimize MSE to fit $rate_{n}^{est}$, with the objective function defined as $\arg\min_{a,b}\frac{1}{n^{{}^{\prime}}}\sum_{n=0}^{n^{{}^{\prime}}}(rate_{n}^{est}-rate_{n})^{2}$, where $n^{{}^{\prime}}$ is the number of recombined layers. The estimation begins after init. Once a new layer is selected, LWS control updates the estimation. Due to the limited updates, the estimation may be inaccurate. To mitigate, we also develop a dynamic search range mechanism that dynamically adjusts the search range in each pool (details in Appendix B.5).
@@ 6 Experiment
Table 1: State-of-the-art (SOTA) baselines.
@@
  TL (Jiang et al., 2022) Single-source A standard transfer learning method by freezing all layers
FT (Oh et al., 2022) Single-source A standard fine-tuning method by updating all layers
ProtoNet (Chen et al., 2019) Single-source A distance-based FSL method
MAML (Chen et al., 2019) Single-source A learning-based FSL method
DAPN (Zhao et al., 2021) Single-source A cross-domain FSL method using adversarial ProtoNet
SemiCMT (Chen et al., 2024) Single-source A cross-modal method using contrastive alignment
GPT2 (Zhou et al., 2023) Single-(LLM) A large language model based method using fine-tuning
MDDA (Zhao et al., 2020) Multi-source A cross-domain FSL method using knowledge distillation
MCW (Lee et al., 2019) Multi-source A cross-domain FSL method using maximal correlation analysis
MetaSense (Gong et al., 2019) Oracle-(GT) An advanced FSL method based on MAML with sufficient target data
@@
miniImageNet (Vinyals et al., 2016) Image - 100 - 3x84x84 Public
miniDomainNet (Zhou et al., 2021) Image - 126 - 3x224x224 Public
Office-31 (Saenko et al., 2010) Image - 31 - 3x224x224 Public
Office-Home (Venkateswara et al., 2017) Image - 65 - 3x224x224 Public
Caltech-101 (Fei-Fei et al., 2004) Image - 101 - 3x224x224 Public
Newsgroup (Thulasidasan et al., 2021) Text - 20 - 100x1x1000 Public
VoxCeleb (Nagrani et al., 2017) Audio - 100 Time-S 1x512x300 Public
OPPORTUNITY (Roggen et al., 2010) IMU 12 11 Time-S 77x1x100 Public
@@
HHAR (Stisen et al., 2015) IMU 9 6 Time-S 6x1x256 Public
WESAD (Schmidt et al., 2018) ECG 15 3 Time-S 10x1x256 Public
@@
ChestX (Wang et al., 2017) Image - 5 - 1x1024x1024 Public
@@
  ResNet18 (Chen et al., 2019) 18 9 3.67 11.18M 2D 3x224x224
ResNet18-1D (Chen et al., 2019) 18 9 0.39 3.9M 1D 100x1x1000
Multi-Conv4 (Zhao et al., 2020) 4x5 4 0.3 1.15M 2D 3x32x32
Multi-ResNet10 (Zhao et al., 2020) 10x5 5 0.7 24.5M 2D 3x84x84
Multi-ResNet18 (Zhao et al., 2020) 18x5 9 18.35 55.9M 2D 3x224x224
Conv4 (Chen et al., 2019) 4 4 0.02 0.07 1D 1x1x500
Conv5 (Gong et al., 2019) 5 5 0.19 30.42M 1D By Input
GPT2 (Zhou et al., 2023) 4x6 6 0.41 82.9M 1D By Input
@@
6.1 Experiment setup
We benchmark XTransfer against SOTA baselines, grouped into single- and multi-source methods (Table 1). Evaluation is conducted on 8 widely-used public datasets under cross-modality FSL settings (Table 2), using their corresponding public pre-trained models as sources, from diverse modalities (e.g., image, text, audio, and sensing) with various backbones such as ResNet18 from PyTorch Hub (Hub, 2022) (Table 3). To evaluate XTransfer, we develop 4 real-world applications and build testbeds on 3 commercial edge devices (e.g., smartwatch, smartphone, Raspberry Pi). Based on these, we collect 4 human sensing datasets (marked as Private in Table 2) as targets, from 40 participants in real-world settings, with ethics approval from the Human Research Ethics Committee of our institute. Additionally, we include 2 public sensing datasets and 1 image dataset as targets for standard benchmarking. We apply standard Leave-One-Out Cross-Validation (LOOCV) (Gong et al., 2019) for evaluating human sensing datasets, and adopt accuracy-to-resource (ATR) ratio, defined as $ATR=Accuracy\mathbin{/}(\alpha Norm(FLOPs))+(1-\alpha)Norm(Params))$ for evaluating resource-accuracy efficiency, where $\alpha$ is set to 0.5 by default (see Appendix A for details).